OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

y-4=-1(x+1) y=-x+3

OpenStudy (anonymous):

y=-1x-5 is slope int form x=y=5=0 is standard

OpenStudy (anonymous):

its not -5 its 3!! go improve your arithmetic

OpenStudy (anonymous):

it can also be x-y=-3 ?

OpenStudy (anonymous):

no y must be positive .... the standard for is Y=mX+c

OpenStudy (anonymous):

m is the gradient and c is the intercept

OpenStudy (anonymous):

oh so wich could it be between this ones?? –x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

x+y=3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

oh all right thank you so much!!

OpenStudy (anonymous):

y must be positive

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

its my pleasure :D

OpenStudy (anonymous):

:D